# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v5

@@ 4 Method
Figure 3: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage (obtains $\mathcal{P}$) and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure 3). We first formulate and analyze the cross-client domain coverage optimization problem. Then, we introduce the FedDCA algorithm, which consists of two main stages: greedy client center selection and client-center-based domain data retrieval. For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder settings, named FedDCA.
@@ 4.1 Optimization Problem
As domain coverage directly affects the in-domain performance of the LLM, FedDCA aims to maximize the domain coverage of the cross-client augmented data $D^{c}$ with respect to the in-domain data distribution $D^{d}$. However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage is unrealistic. Instead, finding a proper client center set $\mathcal{P}$ with optimized domain coverage and using it to perform instruction retrieval on the public data is an approximation problem.
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm Wu (2012). Additionally, in the federated setting, communication cost is always a critical factor. Thus, we formulate the optimization problem as follows:
$\footnotesize\operatorname*{arg\,min}_{\mathcal{P}}$$\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|-\frac{1}{|D^{d}|}\sum_{d\in D^{d}}\left(\max_{p\in\mathcal{P}}\operatorname{sim}(d,p)\right)\right\},$ (4)
s.t. $\quad|\mathcal{P}_{i}|\leq\xi,\quad\forall i\in\{1,2,\dots,N\}$ (5)
The second term is the domain coverage of the selected client center set $\mathcal{P}$, and $sim(\cdot,\cdot)$ is the cosine similarity function. However, this optimization problem is NP-hard.
For simplicity, we let $\xi$ be the constant $N$ (further discussed in Section 6.3). To solve this problem, the time complexity is $\mathcal{O}\left(C_{\xi N}^{N}\left(\mathcal{N}\left(N+N\log N\right)\right)\right)$, where $\mathcal{N}$ is the size of the public data $D^{p}$. As it is a factorial equation, the computational cost increases exponentially with $N$. Furthermore, the enormous amount of public data results in a large $\mathcal{N}$. To enhance computational efficiency, we propose a greedy algorithm as described below.
@@ 4.2 FedDCA
As the goal of FedDCA is to find a proper client center set $\mathcal{P}$ to optimize the objective defined in Eq. 4, we propose a greedy algorithm as shown in Algorithm 1 to solve this problem in polynomial time and obtain a sub-optimal solution. Given the cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ received from each client, which are obtained by clustering local instructions. FedDCA on the server consists of two main steps: greedy client center selection and client-center-based domain data retrieval.
Naive Greedy Client Center Selection.
In this paper’s FedDIT setting, a natural attempt to achieve greedy domain coverage optimization is using the public dataset $D^{p}$ on the server and maximizing $d(D^{d},\mathcal{P})$. Here, each client provides a set of candidate embeddings, and the server greedily picks (or updates) the cluster center to maximize domain coverage on $D^{p}$. While this approach is intuitive, it has three main limitations: 1) High computation complexity when $D^{p}$ is large. Each step of finding the best center $q^{*}$ (to add or update) must re-evaluate coverage w.r.t. $|D^{d}|$ data points, often incurring complexity on the order of $\mathcal{O}\bigl{(}N\xi|D^{d}|(N+N\log N)\bigr{)}$ per selection step. This quickly becomes prohibitively expensive for large $|D^{d}|$. 2) Difficulty in domain coverage computation. The server may have limited knowledge about which domain each public instruction belongs to, or only know it partially. Additionally, the definition of domain is flexible without unified annotation standards - clients’ perceptions of domains may differ from the server’s understanding. This makes it challenging for the server to accurately compute domain coverage. 3) No guaranteed public data. In many realistic FedDIT deployments, the server may not actually have an in-domain public dataset. Clients might rely on external retrieval or self-instruct generation for data augmentation, leaving no $D^{p}$ accessible on the server. The coverage-based greedy approach is then infeasible.
Cluster-center-based Client Center Selection.
To address these issues, we propose an alternative method that does not rely on public dataset. We denote the full set of these uploaded cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ as $\mathcal{C}_{\text{all}}$, which is the proxy for domain coverage space. Since each client must pick one cluster center from $\mathcal{C}_{\text{all}}$. We then define the coverage score for client center selection:
$\quad C(\mathcal{P})$ $=\sum_{z\,\in\,\mathcal{C}_{\text{all}}}\max_{p\in\mathcal{P}}\mathrm{sim}(z,p),$ (6)
$where\quad\mathcal{P}$ =$=\{\,p_{1},\dots,p_{N}\,\},\quad p_{i}\in\mathcal{C}_{\text{all}},$ (7)
and seek to optimize it. This avoids relying on $D^{d}$, greatly reducing coverage-computation overhead and allowing operation without a public dataset.
@@
Specifically, each client $c_{i}$ locally runs $k$-means to produce $\xi$ candidate centers $\mathcal{C}_{i}$, then uploads these centers to the server. The server initializes a selection $\mathcal{P}$ by picking one cluster center from each client (thus $|\mathcal{P}|=N$) and computes the overall coverage solely based on these uploaded centers. In each iteration, the algorithm employs a coordinate-ascent style update: it removes the currently chosen center from one client’s portion of $\mathcal{P}$, tries replacing it with other candidate cluster centers, and keeps the replacement if it improves the coverage; this process is repeated until no further improvement is found (as shown in Algorithm 1).
Algorithm 1 FedDCA: Greedy Client Center Selection
@@
4:$\Omega^{*}\leftarrow C(\mathcal{P})$
5:for $i=1,\dots,N$ do
6:    $\mathcal{P}_{-i}\leftarrow\mathcal{P}\setminus\{p_{i}\}$triangleright Remove $p_{i}$ from the current center set
7:    $\nu^{*}\leftarrow p_{i}$
@@
9:         $C_{i}=C(\mathcal{P}_{-i}\cup\{q\})$
10:         if $C_{i}>\Omega^{*}$ then
11:             $\Omega^{*}\leftarrow C_{i}$; $\nu^{*}\leftarrow q$
@@
14:    $p_{i}\leftarrow\nu^{*}$; $\mathcal{P}\leftarrow\mathcal{P}_{-i}\cup\{p_{i}\}$; $C(\mathcal{P})\leftarrow\Omega^{*}$
15:end for
16:Output: The final center set $\mathcal{P}$ and its coverage $C(\mathcal{P})$.
Since coverage is computed over just $N^{2}$ local centers (i.e., $\mathcal{C}_{\text{all}}$) rather than a large $D^{p}$, the computational overhead is much more manageable. Moreover, by not relying on a public dataset, this approach is applicable for various scenarios, including both data augmentation and data selection.
@@
For each client center $p_{k}$, the server performs dense retrieval (detailed in Appendix A.2) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients. Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
FedDCA: Heterogeneous Encoders.
To enhance client-side computational efficiency and system scalability, we introduce FedDCA. This method employs a small encoder $\omega$ on the client side to reduce computational overhead and a larger encoder $\omega^{\prime}$ on the server side. However, the output dimension of heterogeneous encoders may not be consistent. Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between. We use contrastive learning Khosla et al. (2020) and train $w_{p}$ on the public instructions. For $\forall{x_{i}}\in D^{p}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$. For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$. Let the batch size be $B$ and $\tau$ be the temperature parameter, then the training objective is defined as follows:
$\footnotesize\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (8)
In summary, FedDCA leverages greedy client center selection and domain data retrieval to maximize cross-client domain coverage, leading to better model performance on domain-specific tasks. Additionally, FedDCA employs heterogeneous encoders with server-side feature alignment to reduce client-side computational overhead while maintaining effectiveness.
5 Discussions
Computation & Communication.
Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 4 in the polynomial time. As mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N^{4}\xi^{2}\left(1+\log N\right)\right)$. For domain data retrieval, its time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.

### v6

@@ 5 Method
Figure 2: Overview, a three-stage process. 1) Client-side: Clients $c_{k}$ perform local instructions clustering and send cluster centers $\mathcal{C}_{k}$ to the server. 2) Server-side Augmentation: The server executes diversity-oriented client center selection to maximize domain coverage (obtaining $\mathcal{P}$), retrieves augmented data $D_{k}^{g}$ based on these centers, and distributes it to clients. 3) Collaborative Fine-tuning: Clients collaboratively fine-tune the LLM, exchanging LoRA parameters $\Delta\phi$ with the server.
Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure˜2). We formulate the cross-client domain coverage optimization problem and then introduce the FedDCA algorithm.
@@ 5.1 Optimization Problem
As domain coverage directly affects the in-domain performance of the LLM, FedDCA aims to maximize the domain coverage of the cross-client augmented data $D^{c}$ with respect to the in-domain data distribution $D^{d}$. However, privacy constraints prevent clients from sharing local data, meaning the server lacks direct knowledge of clients’ domain information. Consequently, directly identifying a cross-client dataset that maximizes domain coverage is infeasible. Instead, we leverage client-provided cluster centers ($\xi$ clusters per client, obtained by k-means algorithm Wu (2012)) as privacy-preserving proxies for local data. These centers, typically vector averages, effectively represent local distributions with manageable privacy risks (elaborated in Section˜B.5). The core challenge thus shifts to selecting an optimal set of these client centers, $\mathcal{P}$ (one per client), such that retrieving public instructions based on $\mathcal{P}$ enhances cross-client domain coverage. This formulation provides a tractable approximation to the original optimization problem under inherent privacy constraints.
Additionally, in the FL setting, communication cost is always a critical factor. Thus, we formulate the optimization problem as follows:
$\operatorname*{arg\,min}_{\mathcal{P}}\left\{\eta(\mathcal{P})-d(D^{d},\mathcal{P})\right\},$ (3)
where the first term $\eta(\mathcal{P})=\sum_{i=1}^{N}|\mathcal{P}_{i}|$ represents the communication overhead of transmitting client centers. And the second term is the domain coverage of the selected client center set $\mathcal{P}$, defined in Eq.2.
For simplicity, we let $\xi$ be the constant $N$ (further discussed in Section˜6.3). To better fit the FL environment and enhance the computational efficiency, we propose FedDCA as follows.
@@ 5.2 FedDCA
Optimizing Eq. 3 to find the ideal client center set $\mathcal{P}$ presents significant practical hurdles. First, an exhaustive search over all $C_{\xi N}^{N}$ candidate center combinations is computationally infeasible due to its exponential complexity. Second, directly leveraging a server-hosted public dataset $D^{p}$ to guide client center selection for optimal domain coverage is often impractical. This stems from several factors: 1) privacy constraints limit the server’s knowledge of individual client data distributions and their specific domain focuses; 2) the inherent ambiguity and hierarchical nature of ‘domains’ (e.g., ‘medical’ versus specific medical sub-tasks) complicate precise server-side domain coverage calculation, which is already costly over a potentially massive $D^{p}$; 3) the assumption of a relevant $D^{p}$ being available on the server may not hold, particularly if clients utilize external retrieval or local self-instruct for augmentation.
These challenges underscore the need for a more pragmatic selection strategy. Therefore, we propose FedDCA. As detailed in Alg. 1, FedDCA seeks a sub-optimal solution in polynomial time (discussed in Section˜A.1) and operates on the server through two main stages: 1) a coverage-oriented client center selection mechanism, and 2) client-center-based dense retrieval for data augmentation.
Coverage-oriented Client Center Selection.
To address these challenges, we propose a novel client center selection strategy that, crucially, does not rely on a server-side public dataset $D^{p}$. Let $\mathcal{C}_{\text{all}}=\bigcup_{i=1}^{N}\mathcal{C}_{i}$ be the set of all candidate cluster centers uploaded by the $N$ clients (where each client $c_{i}$ provides $\xi$ centers in $\mathcal{C}_{i}$). Our approach leverages $\mathcal{C}_{\text{all}}$ in two ways: 1) We use $\mathcal{C}_{\text{all}}$ as a proxy for the target domain distribution. Domain coverage is then approximated by evaluating how well a selected subset of centers covers $\mathcal{C}_{\text{all}}$ itself, making the computation tractable and independent of $D^{p}$. 2) Instead of an exhaustive search, we employ an iterative, coordinate ascent-style algorithm (detailed in Alg. 1) to select a final set $\mathcal{P}$ of $N$ client centers. This efficiently yields a high-quality, albeit potentially suboptimal $\mathcal{P}$. Based on the $\mathcal{C}_{\text{all}}$, the domain coverage could be calculated as $d(\mathcal{C}_{\text{all}},\mathcal{P})$.
@@
Specifically, each client $c_{i}$ locally runs $k$-means to produce $\xi$ candidate centers $\mathcal{C}_{i}$, then uploads these centers to the server. The server initializes a selection $\mathcal{P}$ by picking one cluster center from each client (thus $|\mathcal{P}|=N$) and computes the overall coverage solely based on these uploaded centers. In each iteration, FedDCA performs the following update (as shown in Alg. 1): it removes the currently chosen center from $\mathcal{P}$, tries replacing it with other candidate cluster centers in $\mathcal{C}_{\text{all}}$, and keeps the replacement if it improves the cross-client domain coverage; this process is repeated until no further improvement is found.
Algorithm 1 FedDCA: Client Center Selection
@@
[…]
5:while True do
6:  $\mathcal{P}_{-i}\leftarrow\mathcal{P}\setminus\{p_{i}\}$ $\triangleright$ Remove $p_{i}$ from $\mathcal{P}$
7:  $\nu^{*}\leftarrow p_{i}$; $\Omega_{\text{old}}\leftarrow\Omega^{*}$
@@
9:   $\mathrm{C}=d(\mathcal{C}_{\text{all}},\mathcal{P}_{-i}\cup\{q\})$ $\triangleright$ Defined in Eq. 2
10:   if $\mathrm{C}>\Omega^{*}$ then
11:     $\Omega^{*}\leftarrow\mathrm{C};\nu^{*}\leftarrow q$
@@
14:  if $\Omega^{*}=\Omega_{\text{old}}$ then
15:   break
16:  end if
17:  $p_{i}\leftarrow\nu^{*}$; $\mathcal{P}\leftarrow\mathcal{P}_{-i}\cup\{p_{i}\}$; $C(\mathcal{P})\leftarrow\Omega^{*}$; $i\leftarrow(i+1)\bmod N$
18:end while
19:Output: The selected client center set $\mathcal{P}$.
This $D^{p}$-agnostic selection core is computationally tractable and versatile, effective for both data augmentation and targeted data selection through the selected client centers $\mathcal{P}$.
@@
For each client center $p_{k}$, the server performs dense retrieval Zhao et al. (2022) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
In summary, through the coverage-oriented client center selection and domain data retrieval, FedDCA obtains a better cross-client domain coverage of the augmented data $D_{k}^{g}$, which leads to a better model performance on the target domain. For the Discussions part please refer to Section˜A.1.
6 Experiments
To demonstrate the effectiveness of FedDCA, we conduct extensive experiments across various domains and with several baselines. For additional results and analysis, please refer to Appendix˜B.
Method MMLU-Med FPB FiQA TFNS GSM8K
Zero-shot 70.60/- 55.94/- 18.54/- 59.21/- 23.27/-
FedAvg McMahan et al. (2017) 68.40/0.6990 58.25/0.8529 14.18/0.8529 66.62/0.8529 47.46/0.7871
FedProx Li et al. (2020) 69.10/- 56.51/- 14.90/- 66.45/- 47.15/-
SCAFFOLD Karimireddy et al. (2020) 70.20/- 62.71/- 15.27/- 66.49/- 49.27/-
FedAvgM Hsu et al. (2019) 64.70/- 68.14/- 29.27/- 70.32/- 46.85/-
Random Sampling 71.30/0.7940 64.19/0.9196 13.09/0.9196 65.53/0.9196 47.38/0.8651
Direct Retrieval 72.20/0.8830 66.31/0.9293 19.11/0.9293 67.62/0.9293 50.87/0.8967
LESS Xia et al. (2024) 71.00/0.7737 60.56/0.8917 16.00/0.8917 61.14/0.8917 43.13/0.8352
FewFedPIT Zhang et al. (2024c) 68.50/0.8250 56.30/0.8780 14.20/0.8780 59.10/0.8780 42.30/0.8380
Self-Instruct Wang et al. (2022) 71.90/0.8586 59.73/0.9015 20.67/0.9015 66.54/0.9015 50.79/0.8811
KnowledgeSG Wang et al. (2024) 73.13/0.9171 66.00/0.9490 32.72/0.9490 71.10/0.9490 51.20/0.9088
FedDCA+FedAvg 74.50/0.9348 67.24/0.9815 35.27/0.9815 73.32/0.9815 52.46/0.9320
FedDCA+FedProx 72.40/- 72.93/- 38.18/- 77.55/- 51.25/-
FedDCA+SCAFFOLD 73.20/- 72.68/- 33.09/- 75.50/- 50.26/-
FedDCA+FedAvgM 68.90/- 71.45/- 31.45/- 72.52/- 49.76/-
Table 1: Performance (%) and domain coverage of FedDCA and other eleven baselines. Due to limited local data or inappropriate data augmentation strategy, occurs performance degradation in the Unaugmented and Random Sampling settings (underlined values). Different FL strategy does not affect domain coverage. We can see that FedDCA outperforms all other baselines.
6.1 Experimental Setup
Dataset and Evaluation Metrics.
To evaluate the performance of FedDCA, we conduct experiments utilizing a server-hosted public dataset consisting of multiple domains’ data (detailed in Section˜B.1): Alpaca, MedAlpaca Zhang et al. (2023d), FinGPT Yang et al. (2023a), and MathInstruct Toshniwal et al. (2024). Performance is evaluated on MMLU-Med (medical); FPB, FiQA, and TFNS (financial); and GSM8K (math). The former two domains’ metric is Acc. and math’s is Exact Match. We set the number of clients to 10 by default, while each client has 100 local instructions and obtains 1k augmented public instructions by default from the server. Specifically, following the construction of the client’s local data heterogeneity in Section˜4, we set the $\beta$ to 0.1 by default. We also exhibits FedDCA’s strong robustness to different levels of data heterogeneity in Appendix˜C.
Baselines.
We compare FedDCA against a range of baseline methods, categorized into unaugmented and augmented approaches (further details are available in Appendix˜B). Unaugmented Methods. These include Zero-shot inference, where the LLM directly predicts without any fine-tuning, and FedIT. FedIT represents the application of standard federated learning to clients’ local data, for which we employ four widely recognized FL algorithms: FedAvg McMahan et al. (2017), FedProx Li et al. (2020), SCAFFOLD Karimireddy et al. (2020), and FedAvgM Hsu et al. (2019). Augmented Methods. This category encompasses several strategies. Random Sampling serves as a simple baseline, randomly selecting $N_{k}^{p}$ public data instances for each client. Direct Retrieval, a precursor to our method, performs dense retrieval from public data based on each client’s local instructions independently, without optimizing for cross-client domain coverage. LESS Xia et al. (2024) utilizes gradients from a warmed-up LLM on both training and a required validation set for its similarity-based retrieval. For generation-based approaches, Self-Instruct Wang et al. (2022) generates new instructions by prompting a powerful external LLM (e.g., GPT-3.5). FewFedPIT Zhang et al. (2024c) is a federated few-shot method where the global model is used for local in-context learning to synthesize data, aiming to improve both performance and privacy. Finally, KnowledgeSG Wang et al. (2024) employs a server-hosted expert model to generate and refine synthetic data, which clients then use to enhance their local models.
6.2 Main Results
Performance & Domain Coverage.
Table˜1 showcases the performance and corresponding domain coverage of FedDCA against various baselines across three distinct domains. FedDCA consistently outperforms all eleven baselines, demonstrating substantial performance improvements of up to 29.19%. This superior performance is strongly correlated with its ability to achieve the highest domain coverage, surpassing other methods by an average of 4.82% to 21.36%.
While unaugmented methods naturally struggle due to limited local data and thus restricted domain coverage, many existing augmentation strategies also exhibit significant drawbacks. 1) Generation-based methods present a mixed picture: KnowledgeSG, leveraging a powerful server-hosted expert model, achieves strong results while FewFedPIT underperform even unaugmented baselines, potentially due to inconsistent synthetic data quality and diversity. Moreover, these approaches often entail significant resource demands: KnowledgeSG and FewFedPIT involve high computational costs (server-side and client-side), while Self-Instruct incurs monetary expenses for API calls with the privacy leakage. The potential lack of quality and diversity in synthetic instructions further constrains their effectiveness. 2) Random Sampling and Direct Retrieval, while less costly, lack a strategic mechanism to ensure comprehensive cross-client domain coverage. 3) LESS, while performs domain specific retrieval, requiring access to a validation set, and additional computation (warmup training).
In conclusion, FedDCA’s strategic data augmentation substantially enhances domain coverage and model performance, efficiently circumventing the computational burden and inherent quality/redundancy challenges of data generation.
Plug-and-Play.
We demonstrate FedDCA’s plug-and-play capability by combining it with other augmentation methods. After sampling 1k instructions per client through different baselines, FedDCA performs retrieval by the selected client center (200 per client). As shown in Table˜2, despite using only 20% of the sampled data, FedDCA achieves comparable performance with baselines. This highlights FedDCA’s efficient data selection capability based on the cross-client domain coverage, which is further discussed in Section˜B.7.
+ FedDCA MMLU-Med FPB GSM8K
Random Sampling 71.80 (+0.50) 63.92 (-0.27) 47.82 (+0.44)
Direct Retrieval 72.10 (-0.10) 67.78 (+1.47) 52.53 (+1.66)
LESS 70.80 (-0.20) 61.03 (+0.47) 43.92 (+0.79)
Self-Instruct 70.70 (-1.20) 60.21 (+0.48) 50.25 (-0.54)
Table 2: Plug-and-Play. Performance (%) of FedDCA with different data augmentation methods. FedDCA matches baseline’s performance by just 10% of its data through selection.
6.3 Ablation Study
Impact of Different Cluster Number.
The hyperparameter $\xi$ is the number of clusters in the k-means algorithm. The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients. We report the domain coverage of the augmented dataset via FedDCA with different $\xi$ on the three domains in Table˜3. Results show that there is no best $\xi$ for all domains and $\xi=N$ is usually an acceptable choice.
(a) Performance
(b) Domain Coverage
Figure 3: Impact of similarity threshold $\alpha$ on FedDCA’s (a) performance (%) and (b) domain coverage. Results highlight the role of the similarity threshold in enhancing diversity. Lighter bars indicate FedDCA.
Similarity Threshold.
Figure˜3 presents an ablation study comparing FedDCA (w similarity threshold $\alpha$) against FedDCA (w/o $\alpha$). FedDCA achieves superior performance and domain coverage, highlighting the crucial role of the similarity threshold in effectively amplifying the diversity of augmented instructions and subsequently boosting model efficacy. Note that the medical domain remains unchanged, as the similarity of all pairwise embeddings is below 0.7 (the default value for the similarity threshold $\alpha$).
Domain $\xi=10$ $\xi=20$ $\xi=40$ $\xi=80$
Med. 0.9348 0.9478 0.9466 0.9618
Fin. 0.9815 0.9814 0.9819 0.9813
Math. 0.9320 0.9348 0.9344 0.9337
Table 3: Domain Coverage for Different $\xi$ Values. $\xi=N$ is usually an acceptable choice.
We also investigate the impact of different retrieval amounts on FedDCA’s performance, detailed in Section˜B.4.
7 Conclusion
This work advances our understanding of Federated Domain-specific Instruction Tuning (FedDIT), establishing cross-client domain coverage as paramount over data heterogeneity for high performance. FedDCA provides a practical pathway to optimize this coverage, demonstrating significant gains through strategic data augmentation. FedDCA’s plug-and-play capability and validated robustness in challenging scenarios (e.g., scarcity of task-specific public data, large-scale deployments) further positions it as a reliable building block for real-world, privacy-preserving specialized LLM applications. These findings offer both an applicable solution and crucial guidance for future federated system design where domain coverage is the key.
Limitations
While FedDCA provides a robust and scalable framework for federated domain-specific instruction tuning, several avenues remain for further exploration. Although we have empirically demonstrated FedDCA’s resilience to the similar cross-client data distributions (see Appendix˜C), its performance may still be influenced by the quality and diversity of both the server-hosted public data and the client-provided candidate centers. Future work could investigate the adaptive client-side clustering to further enhance coverage and robustness. Additionally, automating hyperparameter selection (e.g., the number of clusters per client) and extending FedDCA to handle evolving data distributions or continual learning scenarios are promising directions to further improve its practicality and generalization in real-world deployments.
Appendix A Methodology
A.1 Discussions
Computation.
The client center selection process, as detailed in Alg. 1, operates on $\mathcal{C}_{\text{all}}$, the set of all $N\cdot\xi$ candidate centers uploaded by $N$ clients (each providing $\xi$ centers). Since FedDCA is an iterative optimization algorithm, let $T$ be the number of passes until convergence. In each pass, the algorithm tries to swap the currently selected center $p_{i}$ with one of the other candidate centers from $\mathcal{C}_{\text{all}}$ ($N\xi-N$ choices). Each potential swap requires re-evaluating the domain coverage $\mathrm{C}$, which is $N\xi(N+N\log N)$. Therefore, the total complexity for one pass is approximately $O(N\xi\cdot(N\xi(N+N\log N)))=O(N^{3}\xi^{2}(1+\log N))$. If the algorithm converges in $T$ passes, the total complexity for client center selection is $O(T\cdot N^{3}\xi^{2}(1+\log N))$. Overall, this complexity is polynomial in the number of clients $N$ ($\xi=N$ by default), rendering the selection process tractable for typical federated learning scenarios where $N$ and $\xi$ are manageable.
To empirically validate the convergence speed ($T$) of our client center selection algorithm, we conducted the client center selection for each domain under the default setup for 100 times. The results show that FedDCA converges remarkably quickly, with an average of 1.1N passes across all experiments. This rapid convergence indicates that the algorithm typically requires only one full iteration through all clients, with occasional minor adjustments in a second round of iteration, making it highly efficient in practice.
Communication.